Jasper Metals is considering installing a new molding machine which is expected to produce operating cash flows of $70,500 per year for 9 years. At the beginning of the project, inventory will decrease by $29,600, accounts receivables will increase by $27,800, and accounts payable will increase by $20,100. At the end of the project, net working capital will return to the level it was prior to undertaking the new project. The initial cost of the molding machine is $300,000. The equipment will be depreciated straight-line to a zero book value over the life of the project. The equipment will be salvaged at the end of the project creating an aftertax cash flow of $82,000. What is the net present value of this project given a required return of 11.7 percent?
Answer #1

Investment in working capital = Increase in Accounts receivables - Decrease in Inventory - Increase in Accounts payables

= 27800-29600-20100 = -$21,900 i.e. release

NPV = Present value of cash inflows - Present value of cash outflows

= -300,000 + 21,900 + 70,500*PVAF(11.7%, 9 years) +82000*PVF(11.7%, 9 years) - 21900*PVF(11.7%, 9 years)

= - 278,100 + 70,500*5.3896 + 60,100*0.3694

= $124,067.74